OpenStudy (anonymous):

Find x when y = 5 in the literal equation 4x + 7y = 51.

jimthompson5910 (jim_thompson5910):

4x + 7y = 51 4x + 7*5 = 51 ... replace y with 5 (since y = 5) 4x + 35 = 51 I'll let you finish up